OpenStudy (anonymous):

( x^2 + x + 1) / (x-1). find the limit as x reaches 1

OpenStudy (anonymous):

are you sure with your equation?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

oh yes, people say it does not exist but i think the answer is 3

OpenStudy (anonymous):

and how did you conclude that?

OpenStudy (anonymous):

using the l'hopital rule

OpenStudy (anonymous):

first get the derivatives then put the limit in.

OpenStudy (anonymous):

i have another question, can you help me with it?

OpenStudy (anonymous):

(x^3 - 7x) / X^3. find limit as x reaches 0. is the answer 1 or does not exist?

OpenStudy (anonymous):

it dose not exist because at the last x you still get over 0.

OpenStudy (anonymous):

did you try long division?

OpenStudy (nikvist):

\[\lim_{x\rightarrow 1}\frac{x^2+x+1}{x-1}=\frac{3}{0}=\infty\]

OpenStudy (anonymous):

okay, thanks

OpenStudy (nikvist):

\[\lim_{x\rightarrow 0}\frac{x^3-7x}{x^3}=\lim_{x\rightarrow 0}\frac{3x^2-7}{3x^2}=\frac{-7}{0}=\infty\]